OpenStudy (meilendurcer):

What is the ratio of 4/5 to 7/15 A. 7/12 B. 1 5/7 C. `11/15 D. 1 4/11

OpenStudy (zzr0ck3r):

(7/15)/(4/5) = (7/15)*(5/4) = (7/3)*(1/4) = ?